# VideoAuteur: Towards Long Narrative Video Generation

### v1

@@ 4.2.2 Learning from Noisy Visual Conditions
To enhance the video generation model’s robustness to imperfect visual embeddings $\mathbf{z}_{t}$ from the auto-regressive director, we fine-tune the model using noisy embeddings $\mathbf{z}_{t}^{\prime}$ defined by:
@@
where $\boldsymbol{\epsilon}\sim\mathcal{N}(0,\sigma^{2}\mathbf{z}_{t})$ represents Gaussian noise, $\mathcal{M}$ is a masking operator that sets a fraction of elements to zero, and $\mathcal{S}$ is a shuffling operator that permutes some embedding dimension. Training with $\mathbf{z}_{t}^{\prime}$ improves the model’s ability to handle noisy visual conditions, improving generation quality and robustness with imperfect embeddings.
@@ 5.1 Experimental Setup
Models. Following Seed-Story [54], we use the SEED-X [13] pretrained 7B multi-modal LLM as our base model and apply LoRA finetuning on our narrative dataset. For video generation, we employ a Sora-like model  [2], which has been pretrained on large-scale video-text pairs and could accept both text and visual conditions.
@@
Data. We use a total of $\sim$32K narrative videos for model developing and also use $\sim$1K videos for validation. All the videos are resized to 448 (short-side) and then center-cropped with 448x448 resolution.
@@
Training & Evaluation. Experiments of interleaved auto-regressive director model are trained with 5,000 steps by default. Training loss is a combination of cosine similarity loss and MSE loss for visual tokens and CrossEntropy loss for language tokens. For visual-conditioned video generation, we use the diffusion loss following DiT [32] and Stable Diffusion 3 [10]. Narrative generation is mostly evaluated on the Youcook2 validation set because of the high-quality of action annotations and the Howto100M validation set is mostly used for data quality evaluation and I2V generation.
5.2 Interleaved Narrative Director
In this section, we discuss three key aspects we explored to improve interleaved auto-regressive model for interleaved narrative visual generation: 1) different visual latent space for visual regression, 2) loss design for accurate latent regression, and 3) the cross-modality regression task.
CLIP beats VAE for interleaved generation. We experiment with three different auto-encoded visual latent spaces for regression: the EMU-2 [42] CLIP-Diffusion autoencoder, the SEED-X CLIP-Diffusion autoencoder, and the KL Variational autoencoder (VAE) used by SDXL. Both SEED-X and EMU-2 use a CLIP vision encoder and a finetuned SDXL diffusion model as the decoder for encoding visual latent. From Figure 5, we can observe that SDXL-VAE achieves the best reconstruction quality. However, in terms of visual generation quality, as shown in Table 5, the CLIP-Diffusion based autoencoders significantly outperform VAE (i.e., +12.2 CLIP-T score and 256.6 better FID). This suggests that CLIP embeddings are more suitable for interleaved visual generation compared to VAE’s latent space. This is reasonable, as SDXL-VAE is not aligned with language and lacks semantic understanding.
Method Autoencoder Style VL Aligned. Recon. Ability CLIP-T FID
SDXL-VAE Variational U-Net High 13.2 286.6
EMU-2 CLIP-Diffusion Medium 25.4 76.7
SEED-X CLIP-Diffusion Low 25.1 30.1
@@ 5.2 Interleaved Narrative Director
Latent scale and direction matters. To determine an effective supervision strategy for visual embeddings, we firstly test the robustness of the latents to pseudo regression errors by rescaling (multiplying by a factor) and adding random Gaussian noise. Figure 6 indicates that both scale and direction are critical in latent regression. Notably, rescaling primarily affects object shape while preserving key semantic information (i.e. object type and location), whereas adding noise drastically impacts reconstruction quality. As shown in Table 6, combining MSE loss (minimizing scale error) and cosine similarity (minimizing direction error) leads to the best generation quality (i.e., +1.9 CLIP score and 3.4 better FID), which further verifies our findings.
Figure 6: Both Scale and Direction Matters. We experiment with pseudo regression errors by altering latent direction and scale using Gaussian noise and scaling factors. The reconstruction results confirm that preserving both scale and direction is essential for accurate latent regression.
SEED-X Latent EMU-2 Latent
Loss Type Training Validation Training Validation
MSE Cos. L2 Dist. Cosine. CLIP FID L2 Dist. Cosine. CLIP FID
0.41 0.82 23.6 31.9 1.3 0.78 25.1 80.1
1.1 0.82 24.1 32.1 2.5 0.79 23.5 115.3
0.41 0.83 25.1 30.1 1.4 0.79 25.4 76.7
@@
Table 6: Regression loss with scale and direction. We track the training convergence and evaluate models with the CLIP-T and FID metrics on the validation set. Both Seed-X and EMU-2 latent space show that a combination of both MSE loss and Cosine Similarity loss considering both scale and direction performs best. SEED-X and EMU-2 original regression loss setting is grayed.
@@
Regression Task L2 Dist. Cosine Sim. CLIP-T FID
@@
Figure 7: Quality Comparison Example. Given a global system prompt—“Step-by-step guide to cooking mapo tofu:”—along with the action, caption, and image embeddings of the first step keyframe, our interleaved director sequentially generates ”actions,” ”captions,” and image embeddings to construct a narrative on how to cook the dish step by step. The first two rows display the directly generated keyframes (decoded from the image latents) using the EMU-2 and SEED-X latent spaces. The generated images are realistic with strong visual consistency but are less aesthetically refined than those produced by state-of-the-art text-to-image models, i.e. SDXL and FLUX.1-s.
Interleaved or language-centric?
We compare the interleaved auto-regressive model, with a language-centric approach that generates images based solely on language states (captions) using models like SD-XL and FLUX.1-schnell. As illustrated in Figure 7, the language-centric method using text-to-image models generates visually appealing keyframes but lacks consistency across frames due to limited mutual information. In contrast, the interleaved generation approach, leveraging language-aligned visual latents, learns a realistic visual style through training. However, it occasionally produces images with repetitive or noisy visual elements, as the auto-regressive model struggles to generate accurate embeddings in a single forward. The human evaluation (Table 8) further validates this point. From Table 8, we can observe that interleaved methods achieve the highest scores in realism (4.3 vs. 3.1) and visual consistency (4.5 vs. 3.4), while language-centric methods achieve the best aesthetic score (4.8 vs. 2.1). We argue that visual consistency is more crucial in long narrative video generation, making interleaved methods preferable.
Latent condition Gen. strategy Aesthetic Realistic Visual consist. Narrative
EMU-2 Latent Interleaved 0.7 1.2 2.9 2.2
SEED-X Latent Interleaved 2.1 4.3 4.5 4.4
Text (SDXL) Language-centric 4.0 2.9 3.3 4.0
Text (FLUX.1-s) Language-centric 4.8 3.1 3.4 4.4
Table 8: Human Evaluation – Interleaved vs. Language-centric. This table compares different latent conditioning and generation strategies based on semantic alignment, aesthetic quality, visual consistency, and narrative coherence. Each aspect is scored with five tiers: 1$\sim$5, score higher is better.
5.3 Visual-Conditioned Video Generation
As detailed in Section 4.2, we fine-tune the model to be directly conditioned on the visual latents generated by our interleaved director, rather than relying solely on initial keyframes. Table 9 compares the keyframe-conditioned approach with our visual embedding-conditioned strategy. Our method improves CLIP-T scores on both validation sets—from 25.9 to 26.4 on YouCook2 and from 26.6 to 27.3 on HowTo100M. Additionally, FVD scores decrease, indicating better video quality ( 557.7 vs. 512.6 on YouCook2, 541.1 vs. 520.7 on HowTo100M). Videos conditioned on visual embeddings demonstrate higher semantic alignment and improved generation quality. Video demos are provided on the demo page and in the supplementary materials.
Visual Condition YouCook2 HowTo100M
CLIP-T FVD CLIP-T FVD
Keyframe 25.9 557.7 26.6 541.1
Embedding 26.4 512.6 27.3 520.7
Table 9: Keyframes vs. Visual Embeddings. Evaluate CLIP-T and FVD scores for video generation conditioned on keyframes versus visual embeddings generated by our interleaved director.
@@ 6 Conclusion
In this paper, we tackle the challenges of generating long-form narrative videos and empirically evaluate its efficacy in the cooking domain. We curate and annotate a large-scale cooking video dataset, capturing clear and high-quality narratives essential for training and evaluation. Our proposed two-stage auto-regressive pipeline, which includes a long narrative director and visual-conditioned video generation, demonstrates promising improvements in semantic consistency and visual fidelity in generated long narrative videos. Through experiments on our dataset, we observe enhancements in spatial and temporal coherence across video sequences. We hope our work can facilitate further research in long narrative video generation.
Supplementary Material
@@
  • Appendix F: Generated Video Examples
@@ Appendix A Data Examples with Annotations
Figures 9 and 11 shows two data examples from our CookGen dataset, annotated with high-quality descriptions that provide detailed multi-modal information of cooking processes. The examples clearly show structured annotations of key actions and corresponding visual descriptions, making the dataset ideal for generating long narrative videos.
@@ Appendix B Additional Data Statistics
Appendix C Data Evaluation Details
C.1 Inverse Video Generation Results
As discussed in Section 3.3.1, high-quality captions, especially with ground truth keyframes, enable effective video reconstruction. We compare ground truth video frames with inversely generated frames using the GT first keyframe and annotated captions, as shown in Figures 14, 15 and 16. The reconstruction aligns well with the narrative, accurately capturing actions, though patterns and interactions differ slightly from the original video. This shows that while the captions convey crucial information for reconstruction, they lack finer visual details, a limitation for current vision-language models and human annotators.
@@ C.1 Inverse Video Generation Results
C.2 Prompt for Video Captioning
@@ C.3 GPT-4o Evaluation on Captions
C.4 Human Evaluation on Captions
@@ Appendix E Action-Caption Matching Pseudo Code
The algorithm processes each video iteratively. For each video, it retrieves all associated actions and their time intervals. Then, for each clip in the video, it calculates the IoU with every action and evaluates the matching conditions. Valid matches, along with their metadata such as clip IDs and descriptions, are stored in a list $\mathcal{M}$. This systematic approach ensures that the matched actions and captions are temporally and semantically consistent, providing high-quality annotations for keyframe visual states.
@@
Appendix F Generated Video Examples
Figures 17 and 18 present two examples of long narrative video generation for cooking “Fried Chicken” and “Shish Kabob,” illustrated step-by-step. The generation process begins with our interleaved auto-regressive director, which generates keyframe visual embeddings and their corresponding captions. These embeddings and captions are then used as conditions for the video generation model, which produces high-quality video clips that effectively narrate the cooking process and emphasize the crucial “action” information. The resulting video clips demonstrate excellent performance in capturing the step-by-step cooking instructions. All video clips are also included in the supplementary materials for further review.

### v2

@@ 4.3.2 Learning from Noisy Visual Conditions
To enhance the robustness over imperfect visual embeddings $\mathbf{z}_{t}$ from the auto-regressive director, we fine-tune the model using noisy embeddings $\mathbf{z}_{t}^{\prime}$ defined by:
@@
where $\boldsymbol{\epsilon}\sim\mathcal{N}(0,\sigma^{2}\mathbf{z}_{t})$ represents Gaussian noise, $\mathcal{M}$ is a masking operator that sets a fraction of elements to zero, and $\mathcal{S}$ is a shuffling operator that permutes the order.
Figure 5: Rolling Context Conditioned Render. We integrate tiled global captions, predicted visual embeddings, and a rolling context of previous keyframes to render new keyframes throughout the narrative. By combining semantic conditioning from textual captions and CLIP embeddings with detailed information from VAE embeddings, the diffusion transformer maintains consistency in visual details such as clothing, food details, and character identities. Generated frames are highlighted with red edges.
Figure 6: Quality comparison on long narrative generation. Here is a case with a narrative topic of “Step-by-step guide to cooking blueberry muffins”. Our interleaved director sequentially generates ”actions,” ”captions,” and image embeddings to construct a narrative on how to cook the dish step by step and then render keyframes. Our method shows state-of-the-art visual quality with superior consistency.
@@ 5.1 Experimental Setup
Models. We initialize the auto-regressive model with [13], a pretrained 7B multi-modal LLM. We initialize the context conditioned render model with FLUX.1 Fill model [27]. For video generation, we employ a video generation model which has been pre-trained on large-scale video-text pairs and could accept both text and visual conditions.
@@
Data. We use a total of $\sim$32K narrative videos for model training and another $\sim$1K videos for validation. All the videos are resized to 448 (short-side) and then center-cropped with 448$\times$448 resolution.
@@
Training & Evaluation. We train the interleaved auto-regressive director model for 5,000 steps by default. Training loss is a combination of cosine similarity loss and MSE loss for visual tokens and CrossEntropy loss for language tokens. For rolling context conditioned render, we use the flow matching loss following FLUX [27]. For visual-conditioned video generation, we use the diffusion loss following DiT [34] and Stable Diffusion 3 [10]. Narrative generation is mostly evaluated on the Youcook2 validation set because of the high-quality of action annotations and the Howto100M validation set is mostly used for data quality evaluation and I2V generation. Please refer to the appendix for implementation details.
Evaluation Metrics. The common metrics CLIP score [17] and FVD [49] are used to assess overall video quality, while the FID [18] score evaluates the quality of the generated keyframes. Additionally, when comparing to state-of-the-art baselines, human evaluation is used to assess generation aesthetics, realism, visual consistency across video clips, and the narrative score which reflects the coherence of the generated cooking steps, and if the cooking process has been successfully completed.
5.2 Rolling Context Conditioning
As detailed in Section 4.2, we leverage the in-context conditioning capabilities of the transformer architecture and adopt a rolling context conditioning strategy to enable DiT to render keyframes with superior visual consistency, while adhering to the extended narrative semantics produced by the interleaved auto-regressive director model. As shown in Figure 5, our keyframe renderer preserves fine visual details and exhibits high visual quality and aesthetics with the help of large-scale pretraining [27]. The reason is that the in-context conditioned VAE features could preserve visual details and the semantics are preserved through the auto-regressive model. Notably, the rolling context conditioning approach allows the renderer to strike a flexible balance between generation efficiency and visual consistency by dynamically adjusting the number of frames generated in each forward pass (i.e., a dynamic number of frames).
5.3 Visual-Conditioned Video Generation
As detailed in Section 4.3, we fine-tune the model to be directly conditioned on the visual latents and generated by our interleaved director and keyframes generated by rolling-context renderer. Table 3 compares the keyframe-conditioned approach with our visual embedding-conditioned strategy. Our method improves CLIP-T [17] scores on both validation sets—from 25.9 to 26.4 on YouCook2 and from 26.6 to 27.3 on HowTo100M. Additionally, FVD scores decrease, indicating better video quality (557.7 vs. 512.6 on YouCook2, 541.1 vs. 520.7 on HowTo100M). Videos conditioned on visual embeddings demonstrate higher semantic alignment and improved generation quality. We also provide qualitative samples on the demo page and in the appendix.
Visual Condition YouCook2 HowTo100M
CLIP-T $\uparrow$ FVD $\downarrow$ CLIP-T $\uparrow$ FVD $\downarrow$
Keyframe 25.9 557.7 26.6 541.1
Embedding (w/o Reg.) 25.5 590.3 26.4 554.3
Embedding (w/ Reg.) 26.4 512.6 27.3 520.7
Table 3: Visual-conditioned Video Generation with Regularization. Evaluate CLIP-T and FVD scores for video generation conditioned on keyframes versus visual embeddings generated by our interleaved director with and without regularization.
5.4 Comparisons of Long Narrative Generation
As most existing narrative generation methods [55, 64] only support image generation, we compare our model with state-of-the-art methods on the task of long narrative keyframe generation. We provide both quantitative comparisons in (§5.4.1) and qualitative comparisons (§5.4.2).
Method Prompting Gen. Metric Human Evaluation
Prompt Src. Cond. CLIP-T FID Aes. Real. Consist. Narr.
SD-XL [35] External Text 27.1 - 4.0 2.9 3.3 N/A
FLUX.1-s [27] External Text 27.9 - 4.8 3.1 3.4 N/A
IC Lora [22] External Text 27.9 34.1 4.7 4.1 4.7 N/A
StoryDiffusion [64] External Text 25.9 36.4 3.9 2.9 3.7 N/A
Vlogger [65] LLM Text 25.5 45.5 4.0 2.4 3.1 3.7
Seed-Story [55] Interleaved V. Emb. 24.1 32.1 1.9 4.1 4.2 4.1
Ours (w.o RCC) Interleaved V. Emb. 26.1 25.3 2.1 4.3 4.5 4.4
Ours (w. RCC) Interleaved T+V. Emb. 28.0 29.4 4.8 4.5 4.8 4.6
Table 4: Quantitative comparisons with metrics and human evaluation. Each method is evaluated by both image generation metrics (CLIP-T and FID) and human ratings. Higher values indicate better performance for all human-evaluation metrics (5 tiers, from 1 to 5, higher is better). SD-XL and FLUX.1-s use narrative captions generated by our model and IC-Lora uses a tiled version. RCC: Rolling Context Conditioning. We use our generated narrative captions for the text-conditioned methods (row 1-5).
5.4.1 Long Narrative Keyframe Generation
We compare our method with leading narrative keyframe generation approaches, including IC Lora [22], StoryDiffusion [64], Vlogger [65], and Seed-Story [55], as well as a language-centric strategy that relies solely on captions (using models such as SD-XL [35] and FLUX.1-schnell [27]). Except for IC Lora and Seed-Story, which are fine-tuned on our CookGen dataset for two epochs, all other methods follow their official inference guidelines with the official checkpoints. As shown in Table 4, our approach achieves the highest generation scores, with a CLIP-Text score of 28.0 and an FID score of 25.3. We also conduct a human evaluation (Table 4) using a five-tier rating scale, where higher is better. Our method attains top performance in aesthetics (4.8 vs. 4.7, IC Lora), realism (4.5 vs. 4.1, Seed-Story), and visual consistency (4.8 vs. 4.7, IC Lora), as well as the highest narrative score of 4.6. These results demonstrate that our method achieves state-of-the-art performance for long narrative generation.
5.4.2 Qualitative Comparisons
In Figure 6, we compare our method with state-of-the-art long narrative keyframe generation approaches, including StoryDiffusion, Vlogger, and Seed-Story, and observe that our results maintain superior visual quality and consistency. In particular, our keyframes balance realism with appealing aesthetics while preserving character identities and smooth transitions. In contrast, competing methods often exhibit color inconsistencies or lose track of concepts—Vlogger occasionally produces uneven color schemes between frames, StoryDiffusion can introduce visual confusion, and Seed-Story sometimes generates mismatched clothing across different scenes. This comparison aligns with the human evaluation results in Table 4, demonstrating our method achieves state-of-the-art performance for long narrative visual generation. The generated keyframes can be extended into full video clips with consistent visuals and coherent storytelling.
5.5 Ablation Studies
In this section, we ablate important designs in VideoAuteur, which improve the interleaved auto-regressive model and the visual-conditioned video generation model for interleaved narrative visual generation.
@@ 5.5 Ablation Studies
Latent scale and direction matter. To determine an effective supervision strategy for visual embeddings, we firstly test the robustness of the latents to pseudo regression errors by rescaling (multiplying by a factor) and adding random Gaussian noise. Figure 19 indicates that both scale and direction are critical in latent regression. Notably, rescaling primarily affects object shape while preserving key semantic information (i.e. object type and location), whereas adding noise drastically impacts reconstruction quality. As shown in Table 5, combining MSE loss (for scale ) and cosine similarity (for direction) leads to the best generation quality, improving CLIP-T by 1.5 points and reducing FID by 1.8 points compared to using MSE alone.
Loss Type Training Validation
MSE Cosine L2 Dist $\downarrow$ Cosine Sim. $\uparrow$ CLIP-T $\uparrow$ FID $\downarrow$
0.41 0.82 23.6 31.9
1.1 0.82 24.1 32.1
0.41 0.83 25.1 30.1
@@
Table 5: Both scale and direction matter. We track the training convergence and evaluate models with the CLIP-T and FID metrics on the validation set. The combination of both MSE loss and Cosine Similarity loss performs best on the validation metrics.
@@
Regression Task L2 Dist $\downarrow$ Cosine Sim. $\uparrow$ CLIP-T $\uparrow$ FID $\downarrow$
@@
Learn from noisy visual conditions. Table 7 presents an ablation study examining the effect of robustness regularization on the visual-conditioned video generation model. We evaluate the generated videos using CLIP-T and FVD. The progressively improved results from 26.4 to 27.3 on CLIP-T and 554.3 to 520.7 on FVD demonstrate the effectiveness of our regularization strategy, which combines random masking, Gaussian noise, and shuffling.
Regularization Setting CLIP-T $\uparrow$ FVD $\downarrow$
Naive Baseline 26.4 554.3
+Random Masking 26.9 539.7
+Random Gaussian. Noise 27.2 522.1
+Random Shuffling 27.3 520.7
Table 7: Learn from Noisy Visual Conditions. Our training regularization strategy enhances the robustness of the visual-conditioned video generation model. Specifically, we apply random masking and shuffling at a rate of 25%, and introduce Gaussian noise with 0.5 std of the embeddings of two thousand samples.
@@ 6 Conclusion
In this paper, we tackle the challenges of generating long-form narrative videos and empirically evaluate its efficacy in the cooking domain. We curate and annotate a large-scale cooking video dataset, capturing clear and high-quality narratives essential for training and evaluation. Our proposed two-stage auto-regressive pipeline, which includes a long narrative director, a rolling context conditioned keyframe renderer and a visual-conditioned video generation model, demonstrates promising improvements in semantic and visual consistency in generated long narrative videos with an unified pipeline. Through experiments on our dataset, we observe enhancements in spatial and temporal coherence across video sequences. We hope our work can facilitate further research in long narrative video generation.
Appendix
@@
  • Appendix F: CLIP beats VAE for interleaved generation
    • Introduces three autoencoders (EMU-2, SEED-X, SDXL-VAE) and compares reconstruction vs. generation performance.
    • Demonstrates that CLIP-diffusion embeddings (EMU-2, SEED-X) outperform SDXL-VAE in language-driven visual generation due to better vision-language alignment.
  • Appendix G: Generated Video Examples
@@ Appendix A Data Examples with Annotations
Figures 8 and 10 shows two data examples from our CookGen dataset, annotated with high-quality descriptions that provide detailed multi-modal information of cooking processes. The examples clearly show structured annotations of key actions and corresponding visual descriptions, making the dataset ideal for generating long narrative videos.
@@ Appendix B Additional Data Statistics
Appendix C Annotation Quality Reverification Details
High-quality captions are essential for narrative visual generation. To verify the quality of our annotations, we build an evaluation pipline of inverse generation (§C.1) and visual understanding through VLM experts (§C.2).
C.1 Inverse Video Generation
This evaluation is motivated by the understanding that high quality captions, when combined with ground truth keyframes, more effectively reconstruct the original videos. We evaluate the dataset’s ability to reconstruct original videos using the annotated captions, with and without conditioning with ground truth keyframes. For this evaluation, we assess the validation set ($\sim$5,000 video clips). We measure reconstruction quality using FVD [49]. The results, shown in Table 14, indicate that our captions capture sufficient semantic information, enabling effective representation of the original videos. When generating with ground-truth keyframes, the video quality is very high and closely aligned with the original videos, as shown by the low FVD score (116.3). Without keyframes, the captions alone still provide reasonable alignment. Examples of reconstructed videos are included in the supplementary materials.
Figure 13: Annotation quality evaluation pipeline. We verify our annotation quality through a pipeline of two major aspects: 1) inverse video generation 2) GPT-4o and human evaluation.
Validation Set w/. GT keyframe W/o. GT keyframe
# Clips FVD FVD
5504 116.3 561.1
Figure 14: High-quality captions enable inverse video generation. We utilize the annotated captions and actions to inversely generate video clips using a pretrained Text-to-Video diffusion model. Higher reconstruction fidelity (i.e., similarity to the original videos) indicates superior captions and actions. Our inversely generated videos achieve very low FVD scores compared to the original videos, highlighting the high quality of our annotations.
GPT-4o Evaluation Human Evaluation
Score (0-100) Qwen2-VL-72B Ours Qwen2-VL-72B Ours
98.0 95.2 79.3 82.0
C.2 Semantic Consistency across VLM Experts
C.3 Inverse Video Generation Results
As discussed in Section C.1, high-quality captions, especially with ground truth keyframes, enable effective video reconstruction. We compare ground truth video frames with inversely generated frames using the GT first keyframe and annotated captions, as shown in Figures 15, 16 and 17. The reconstruction aligns well with the narrative, accurately capturing actions, though patterns and interactions differ slightly from the original video. This shows that while the captions convey crucial information for reconstruction, they lack finer visual details, a limitation for current vision-language models and human annotators.
@@ C.3 Inverse Video Generation Results
C.4 Prompt for Video Captioning
@@ C.5 GPT-4o Evaluation on Captions
C.6 Human Evaluation on Captions
@@ Appendix D Implementation Details
Image Resolutions $448\times 448\times 4$
Optimizer AdamW
Optimizer hyperparameters $\beta_{1}=0.9,\beta_{2}=0.999,\epsilon=10^{-8}$
Peak learning rate $5\times 10^{-4}$
Learning rate schedule Linear warm-up, constant
Gradient clip $1.0$
Total training steps $5,000$
Warm-up steps $500$
Batch size $16$
Optional Masking Probability 0.25
Numerical precision bfloat16
Table 11: Implementation details of the rolling context conditioned render.
Configuration Setting
@@ Appendix E Action-Caption Matching Pseudo Code
The algorithm processes each video iteratively. For each video, it retrieves all associated actions and their time intervals. Then, for each clip in the video, it calculates the IoU with every action and evaluates the matching conditions. Valid matches, along with their metadata (clip info and descriptions), are stored in a list $\mathcal{M}$. This systematic approach ensures that the matched actions and captions are temporally consistent, providing high-quality annotations for keyframe visual states.
@@
Appendix F CLIP beats VAE for interleaved generation.
We experiment with three different auto-encoded visual latent spaces for regression: the EMU-2 [44] CLIP-Diffusion autoencoder, the SEED-X CLIP-Diffusion autoencoder, and the KL Variational autoencoder (VAE) used by SDXL. Both SEED-X and EMU-2 use a CLIP vision encoder and a finetuned SDXL diffusion model as the decoder for encoding visual latent. From appendix Figure 18, we observe that SDXL-VAE achieves the best reconstruction quality. However, in terms of visual generation quality, as shown in Table 13, the CLIP-Diffusion based autoencoders significantly outperform VAE (i.e., +12.2 CLIP-T score and 256.6 better FID). This suggests that CLIP embeddings are more suitable for interleaved visual generation compared to VAE’s latent space. This is reasonable, as SDXL-VAE is not aligned with language and lacks semantics.
Method Autoencoder Style VL Aligned. Recon. Ability CLIP-T FID
SDXL-VAE Variational U-Net High 13.2 286.6
EMU-2 CLIP-Diffusion Medium 25.4 76.7
SEED-X CLIP-Diffusion Low 25.1 30.1
Figure 19: Both Scale and Direction Matters. We experiment with pseudo regression errors by altering latent direction and scale using Gaussian noise and scaling factors. Reconstruction results confirm that preserving both scale and direction is important for latent regression.
Appendix G Generated Video Examples
Figures 20 and 21 present two examples of long narrative video generation for cooking “Fried Chicken” and “Shish Kabob,” illustrated step-by-step. The generation process begins with our interleaved auto-regressive director, which generates keyframe visual embeddings and their corresponding captions. These embeddings and captions are then used as conditions for the video generation model, which produces high-quality video clips that effectively narrate the cooking process and emphasize the crucial “action” information. The resulting video clips demonstrate excellent performance in capturing the step-by-step cooking instructions. All video clips are also included in the supplementary materials for further review.